OpenStudy (anonymous):

I need serious help please...been 8 and 1/2 hours trying to figure out this probability... 126C20/300C20 (rounded to six decimal places as needed). Can anyone help me please?

Directrix (directrix):

C(126,20) = 8.506107 * 10 ^22 C(300,20) = 7.500433 * 10^30 (8.506107 * 10 ^22) / (7.500433 * 10^30) = 1.134082* 10 ^ (-8) = C(126,20)/C(300,20). approx or .00000001134082.